# You Only Click Once: Single Point Weakly Supervised 3D Instance Segmentation for Autonomous Driving

### v1

@@ 2 Related Work
Weakly Supervised 3D Instance Segmentation. Point cloud segmentation has made significant progress in fully supervised settings. However, dense point-wise annotation is costly. To reduce the annotation burden, some work [35, 11, 45, 4, 32, 33, 17, 24, 43, 12] has explored using various weak supervision signals.
@@ 4.4 Ablation Study and Analysis
Table 3: Ablation Study for VFM-PLG Module.
@@
Fine-tuning with the YoCo. Following  [12], we fine-tune the trained network within our YoCo framework across different proportions of fully supervised data. As shown in Figure 1, fine-tuning with a mere 0.8% of the fully supervised data achieves performance comparable to that obtained with full supervision. (59.27% vs. 59.26%). Moreover, when increasing the fully supervised data utilization to 5%, our approach surpasses the fully supervised performance of the state-of-the-art PTv3.
@@
Table 7: Generality Experiment for the YoCo.
@@ 6 Appendix
  • Pseudo Label Quality Comparison. In Section 6.1 we provide a comparison of the quality of pseudo labels generated by different methods.
@@
6.1 Pseudo Label Quality Comparison
Figure 5 demonstrates the effectiveness of our method in generating high-quality pseudo labels, and our VFM-PLG module achieves nearly a 10% performance improvement compared to the method that uses only clicks as prompts.
6.2 TSU Algorithm Implementation
@@
of points whose current voxelization coordinates a- re $(v_{x_{a}},$ $v_{y_{a}},v_{z_{a}})$, and each set contains the coord- inates $(x_{a},y_{a},z_{a})$ and confidence scores $s_{a}$ for the current voxel $v_{a}$.
@@
             $S\left[v_{x_{a}},v_{y_{a}},v_{z_{a}}\right]=\arg\max\left(\bar{s}\right)$;

### v2

@@ 2 Related Work
Weakly Supervised 3D Instance Segmentation. Point cloud segmentation has made significant progress in fully supervised settings. However, dense point-wise annotation is costly. To reduce the annotation burden, some work [36, 12, 46, 4, 33, 34, 18, 25, 44, 13, 10] has explored using various weak supervision signals.
@@ 4.4 Ablation Study and Analysis
Table 3: Ablation Study of VFM-PLG Module.
@@
Fine-tuning with the YoCo. Following  [13], we fine-tune the trained network within our YoCo framework across different proportions of fully supervised data. As shown in Figure 1, fine-tuning with a mere 0.8% of the fully supervised data achieves performance comparable to that obtained with full supervision (59.27% vs. 59.26%). Moreover, when increasing the fully supervised data utilization to 5%, our approach surpasses the fully supervised performance of the state-of-the-art PTv3.
@@
Table 7: Generality Experiment of the YoCo.
@@ 6 Appendix
  • More Experiments. In Section 6.2, we discuss the impact of YoCo on the field of autonomous driving.
  • Visualization. In Section 6.3, we provide some visualization results.
6.1 Algorithm Implementation
@@
of $n$ points whose current voxelization coordinates are $(v_{x_{a}},$ $v_{y_{a}},v_{z_{a}})$, and each set contains the coord- inates $(x_{a},y_{a},z_{a})$ and confidence scores $s_{a}$ for the current voxel $v_{a}$.
@@
             $S\left[v_{x_{a}},v_{y_{a}},v_{z_{a}}\right]=\arg\max_{cls}\left(\bar{s}\right)$;
@@
6.2 More Experiments
Pseudo label quality comparison. Figure 5 demonstrates the effectiveness of our method in generating high-quality pseudo labels, and our VFM-PLG module achieves nearly a 10% performance improvement compared to the method that uses only clicks as prompts.
Results on nuScenes. We conduct experiments on the nuScenes dataset, using only the point clouds projected onto the image. Data augmentation for single-frame data included global rotation, translation, scaling, random flipping, and shuffling. As shown in Table 8, our method YoCo achieves a 14% improvement in mAP and a 17.868% improvement in mIoU compared to the Click-based method, demonstrating its effectiveness.
Sup. Method mAP mIoU Full SparseUnet 49.56 68.056 Cylinder3D 45.74 66.198 PTv3 46.1 71.972 Weak Click 19.35 37.686 YoCo (Ours) 33.35 55.554
Table 8: Performance Comparisons on NuScenes Validation Dataset.
Greater click range error. To evaluate the robustness of YoCo for larger target sizes in autonomous driving, we conduct additional experiments with error ranges of 1m, 1.5m, and 2m for vehicle classes, as shown in Table 9. Expanding the click range to 2.0m results in stable performance and even improvement, demonstrating the robustness of our method. These experiments suggest that the BEV center is not the optimal click, and identifying the optimal click will be explored in future work.
Range 0.0 0.1 0.5 1.0 1.5 2.0 AP 67.69 67.28 67.70 67.22 67.60 68.06 IoU 81.136 82.288 82.229 84.350 85.400 83.885
Table 9: Greater Click Range Error for Vehicle.
False positives (FP) click. We consider that some small objects (such as Construction Cone/Pole) may be mislabeled as people in the annotation. To address this issue, we propose incorporating 2D semantic information to improve segmentation accuracy. Specifically, we intentionally introduce Construction Cone/Pole objects as Pedestrian noise in each scene to simulate real-world mislabeling scenarios. As demonstrated in Tab. 10, our method achieves 53.48% mAP and 74.936% mIoU when using ground truth (GT) boxes, which is comparable to the performance of YoCo. Furthermore, even with imprecise predictions from YOLO, our approach maintains robust performance, validating the efficacy of leveraging 2D semantic information for verification.
Box GT YOLO mAP 55.12 53.48 mIoU 75.120 74.936
Table 10: YoCo with FP Problem.
Discard click analysis. To verify whether all foreground objects need to be clicked, we conduct a random discard experiment. It is noteworthy that our default method inherently discards objects with insufficient point cloud representations in the Bird’s Eye View (BEV) perspective, resulting in a 16.31% object discard rate. We further evaluate the impact of additional discarding by randomly removing instances with probabilities of 10%, 20%, and 30%, as detailed in Table 11 (b), (c), and (d). The experimental results indicate that as the discard rate increases, the network performance gradually declines. However, even with a discard proportion of 37.19%, our method still achieves 53.96% mAP and 73.248% mIoU, demonstrating the robustness of YoCo.
Method Proportion mAP mIoU (a) YoCo 16.31% 55.35 74.770 (b) Discard Object 24.65% (10%) 54.29 73.742 (c) 37.19% (25%) 53.96 73.248 (d) 58.10% (50%) 46.68 70.008
Table 11: Ablation Study of Drop Click.
Scalability across sequences. In Table 12, we assess the efficacy of YoCo under limited annotation scenes by subsampling point cloud sequences at intervals. Specifically, in Table 12 (b) and (c), we report performance metrics using 1/3 and 1/2 of the annotated sequences, respectively. Notably, even with only 1/3 of the annotation data, our method achieves 53.13% mAP and 72.304% mIoU, highlighting its scalability and robustness in resource-constrained settings.
Interval mAP mIoU (a) 1 55.35 74.770 (b) 2 53.90 72.999 (c) 3 53.13 72.304
Table 12: Ablation Study of Across Sequences.
Unsupervised application of YoCo. To extend the applicability of our approach, we investigate the use of the 2D detection model YOLO for unsupervised point cloud segmentation tasks. Specifically, we align YOLO’s detection categories with the annotation categories of the Waymo dataset. Since YOLO does not directly support the Cyclist category, we propose a heuristic strategy: the Cyclist category is inferred by combining detections from the intersection of the Person and Bicycle categories. As shown in Table 13, our unsupervised approach achieves significant performance improvements, with a 5.59% increase in mAP and a 4.672% increase in mIoU compared to the click-based method. These results validate the effectiveness of leveraging 2D detection models for unsupervised point cloud segmentation tasks.
Supervision Annotation Model mAP mIoU Full 3D Mask SparseUnet 59.26 79.505 Weak Click SparseUnet 40.19 67.510 Click YoCo 55.35 74.770 Unsupervised YOLO YoCo 45.78 72.182
Table 13: Performance Comparison of Supervision Strategies on the Waymo Validation Dataset. indicates pseudo labels generated by SAM using the corresponding click annotations as prompts, represents pseudo labels produced by our proposed VFM-PLG module, and YOLO refers to pseudo labels derived from YOLO prediction results.
6.3 Visualization
In Figure 6, we evaluate the performance of SparseUnet trained with various annotation strategies. Our proposed method exhibits superior visual quality, significantly outperforming click-based approaches by generating more accurate and spatially coherent predictions. This improvement is largely driven by the integration of deep prior information into the pseudo label generation process, which enhances the quality of the generated labels.
In Figure 7, we provide visualization segmentation results for some scenes.
Figure 6: Visiualization for Cyclist.
Figure 7: Visiualization for Scenes.